A market research firm supplies manufacturers with estimates of the retail sales of their products from samples of retail stores. Marketing managers are prone to look at the estimate and ignore sampling error. An SRS of 29 stores this year shows mean sales of 77 units of a small appliance, with a standard deviation of 15 units. During the same point in time last year, an SRS of 18 stores had mean sales of 92.75 units, with standard deviation 12.1 units. A decrease from 92.75 92.75 to 77 is a drop of about 20% .

Construct a 95% confidence interval estimate of the difference μ1−μ2 μ 1 − μ 2 , where μ1 μ 1 is the mean of this year's sales and μ2 μ 2 is the mean of last year's sales.

the margin of error is ? ( 8.43642) is the answer I  got and it says its close to the correct response
Answer #1

Ans:

df=29+18-2=45

critical t value,t*=tinv(0.05,45)=2.0141

pooled standard deviation=SQRT(((29-1)*15^2+(18-1)*12.1^2)/(29+18-2))=13.97535

standard error for difference=13.97535*SQRT((1/29)+(1/18))=4.1935

Margin of error=2.0141*4.1935=8.44613

Point estimate=77-92.75=-15.75

95% confidence interval estimate of the difference

=-15.75+/-8.45

=(-24.20, -7.30)